HUBBLE SPACE TELESCOPE DAILY REPORT #2250 PERIOD COVERED: 0000Z (UTC) 11/12/98 - 0000Z (UTC) 11/13/98 Daily Status Report as of 317/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Seven Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. As described in HSTAR 6768 and yesterday's report, the acquisition for the third iteration of this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as planned, and no further anomalies were noted. 1.4 Completed Two Sets of NIC/1/2 7243 (High Resolution Imaging of Io) The Near Infrared Camera (Cameras One and Two) was used to make observations of Jupiter's moon Io. The technique of super-resolution imaging will be used to obtain near-IR images of Io with sufficient resolution to map the surface on this volcanically active satellite of Jupiter. Observations were also made of standard star SAO146868. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed WF/PC-2/STIS/CCD/NIC/1/2/3 7307 (The Formation and Evolution of Rich Star Clusters in the LMC) The WF/PC-2, the Space Telescope Imaging Spectrograph (CCD), and the Near Infrared Camera (Cameras One, Two and Three) were used to make observations of the rich star clusters which are located within the Large Magellanic Cloud in order to study their formation and evolution. These observations will help us understand the origin and evolution of rich star clusters in our own galaxy. The observations were completed as scheduled, and no anomalies were reported. 1.6 Completed STIS/CCD/MA1 7437 (Understanding High-Redshift and Starburst Galaxies: A UV Spectroscopic Survey of O-Stars in the SMC) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make ultraviolet spectrographic observations of galaxy NGC-1868-CE. The data from this survey, of importance in many areas of astrophysics, will immediately be made public as no proprietary period is requested. A wavelength calibration image was also taken. As described in HSTAR 6768 and yesterday's report, the acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.7 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. As described in HSTAR 6768 and yesterday's report, the acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as planned, and no further problems were reported. 1.8 Completed Four Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. As described in HSTAR 6768 and yesterday's report, the acquisition for the first two iterations of this proposal defaulted to fine lock backup on one FGS only. Otherwise, the proposal completed with no further reported problems. 1.9 Completed WF/PC-2 8059 (POMS Test Proposal: WF-2 Targeted Parallel Archive Proposal) The WF/PC-2 was used to make observations in parallel with the prime science observations. The parallel opportunities available with WF/PC-2 in the neighborhood of bright galaxies are treated in a slightly different way from the normal pure parallels. Local group galaxies offer the opportunity for a closer look at young stellar populations. As described in HSTAR 6768 and yesterday's report, the acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.10 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.11 Completed WF/PC-2 6458 (Imaging of QSO Host Galaxies at Z >2) The WF/PC-2 was used to make observations of the QSO/host galaxy 0225-014. QSOs of redshift larger than 2 have been resolved with ~arcsecond resolution on 4m-class ground-based telescopes. Both line and continuum flux have been detected over a few arcseconds, but the nature of the extended emission, and its relation to the nuclear energy source is unclear. As described in HSTAR 6768 and yesterday's report, the acquisition for this proposal defaulted to fine lock backup on one FGS only. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.12 Completed Seven Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed Two Sets of NIC/2/3 8063 (NICMOS Pure Parallel Sponge - Backup) The Near Infrared Camera (Cameras Two and Three) was used to make observations taken in parallel with the prime science observations. This proposal is PRE-the new FOM default. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed NIC/2 7233 (Dust Disks Around Main Sequence Stars) The Near Infrared Camera (Camera Two) was used to make observations of the dust disk around main sequence stars. The observations were made with the coronagraph to minimize the effects of glare from the bright central star. A dark calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.15 Completed NIC/2 7226 (Search for Massive Jupiters) The Near Infrared Camera (Camera Two) was used to make a corographic search for massive planets around nearby, young main sequence stars. Because of the extreme youth of these objects, any low-mass brown dwarf and planetary companions will still be in a higher luminosity phase and thus easily detectable. A dark calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed NIC/1/2/3 7227 (A Search for Low Mass/Sub-Luminous Companions to M-Stars) The Near Infrared Camera (Cameras One, Two and Three) was used to make a search for low mass/sub-luminous companions to low mass M class stars. Knowledge of stellar and sub-stellar masses and luminosities at and below the ~0.08 M{sun} H-burning limit is of fundamental importance. A dark calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.17 Completed WF/PC-2 7377 (Dwarf Elliptical Galaxy Snapshot Survey II) The WF/PC-2 was used to make observations of the dwarf elliptical galaxy VCC896 as part of Snapshot Survey II. This survey will provide important information on the globular cluster systems of such galaxies and the properties of their cores and nuclei. The observations were completed as planned, and no anomalies were reported. 1.18 Completed NIC/1/2 7329 (The Nature of the Damped LyAlpha Absorbers -- A New Study of Young Galaxies) The Near Infrared Camera (Cameras One and Two) was used to make observations of the young galaxy Q1144-0723 in order to study the nature of its damped Lyman-Alpha Absorbers (DLAs). The DLA line systems, occasionally found in quasar spectra, provide one of only a few practical methods of detecting and studying significant numbers of galaxies at high redshifts. The observations were executed as scheduled, and no anomalies were noted. 1.19 Completed Two Sets of WF/PC-2 6673 (Black Holes and Gas Disks in a Complete Sample of Radio Loud UGC Ellipticals) The WF/PC-2 was used to make observations of radio load, elliptical galaxy UGC08433 in order to search for possible black holes and gas disks. Hydrogen-Alpha and broad band observations were taken to determine the luminosity profiles and stellar populations (on the tens of parsec scale), and detect nuclear regions or disks of gas and dust. The observations were completed as planned, and no anomalies were reported. 1.20 Completed NIC/2 7897 (Probing Planetary Formation Around Main-Sequence Stars: A Snapshot Survey) The Near Infrared Camera (Camera Two) was used to make snapshort survey images of the main sequence star HR2483 probing for planetary formation. The IR excess main-sequence stars host cool circumstellar dust disks, which are most likely replenished by the debris of unseen planetesimals. The observations were completed as scheduled, and no problems were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 8 Successful : 8 Scheduled Reacquisitions : 8 Successful : 8 2.2 FHST Updates: Scheduled : 18 Successful : 18 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared twice. A TTR was written to document IP data hits at 316/0934Z and at 316/2320Z. A MA/GN mode SSR Dump test was successfully supported at 316/1758Z. As documented in HSTAR 6769, a NICMOS EDAC single bit error occurred at 316/143055Z. The counter incremented to 42 and then 43. The ground system limit was adjusted. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH